OpenStudy (lxelle):

company makes a rights issue of 150,000 shares of $1 at a price of $1.50. how to record rights issue?

OpenStudy (cj49):

Bank a/c Share Capital a/c Share Premium a/c

OpenStudy (lxelle):

Hi do you count those?

OpenStudy (anonymous):

Wouldn't you just multiply 150, 000 by 1.5?

OpenStudy (lxelle):

Are you sure?

OpenStudy (cj49):

do u need the ledger entry

OpenStudy (lxelle):

No.

OpenStudy (anonymous):

Well if the 150, 000 shares are $1 each and they make them $1.50, then it would be 150, 000 shares of $1.50, or 150, 000 X 1.5

OpenStudy (lxelle):

https://thol.sunway.edu.my/examdbase/alv/acc/p1/acc_p1_j08.pdf#_=_ What about this? For question 19

OpenStudy (lxelle):

Mind helping me out with the answer?

OpenStudy (lxelle):

@cj49 help please?

OpenStudy (cj49):

i guess its A m not sure tho..

OpenStudy (lxelle):

Yeah you're right. Can you tell me how do you get that?

OpenStudy (cj49):

do u now what does Share Premium and Revenue reserves mean

OpenStudy (lxelle):

Yeah . I do I just don't know how to work the figures out.

OpenStudy (cj49):

200+80+160=440 they issued more 1000000 shares of 1$ each at 1.40 per share=140 440+140=580 the actual price was 1$ but they issued it at 1.40$ therefore 40 would be the share premium 580-40=540 if ull look at the options then ull figure it out the A fits the situation

OpenStudy (lxelle):

Why is it $500000 ordinary share capital? How did you split it?

OpenStudy (cj49):

the total sum is 580 out of that 40 wud be share perimium so we are now left with 540 so option B and C wud get eliminated bc in B the share premium is 80 and in C its nil and in option D the total would sum up to 620 i.e540+40+40=620...thats why A had to be the right ans

OpenStudy (lxelle):

I see there thanks man you're really good at explaining. :) don't mind helping me out with question 30? https://thol.sunway.edu.my/examdbase/alv/acc/p1/acc_p1_j08.pdf#_=_

OpenStudy (cj49):

i guess its B

OpenStudy (lxelle):

Yes bingo! How?

OpenStudy (cj49):

lol....1000x12=12000 2000x4=8000 total 20000 now they have asked what does the company should do to increase the PROFIT by 10000=20000+10000=30000 i guess u need to solve this 30 ques in 60 min so they wnt give u problems which wud consume more than 2min. in option A 30% increase in X wud gve=1300x12 u dun need to calculate this bc its just an increase of 300x12=3600 But in option B its 500x12=6000 and 1000x4=4000=10000

OpenStudy (lxelle):

Ahh, okay. JUSt realize that my lecturer haven't touched this topic yet. Anyways thanks! You're really good in accountings. By any chance are you doing ACCA?

OpenStudy (cj49):

thanks..mt dad is a Chartered Accountant so i Audit reports and other stuffs foh him

OpenStudy (cj49):

my*

OpenStudy (lxelle):

I see. :)

OpenStudy (cj49):

c:

OpenStudy (lxelle):

I think I might have more accounting questions to clarify and all I'm not sure if you could help me out with those next time? Hope that it's not too troublesome

OpenStudy (cj49):

lol nopes... do tag me ill be der...

OpenStudy (lxelle):

Thankss! I'm really grateful! Haha. It's really difficult to find an accounting student here.

OpenStudy (cj49):

anytime...lol..yeah if m offline then just pm the problem ill revert back when ill see it

OpenStudy (lxelle):

Alright. :)

OpenStudy (cj49):

c: